OpenStudy (anonymous):

8x-5y=5 3x-8=-y What is the solution in an ordered pair?

OpenStudy (anonymous):

3x-8=-y rewrite is: 3x+y=8 (-8)*(3x+y) =8 -24x-8y=-64 (3)*(8x-5y) =5 24x -15y=15 -24x-8y=-64 24x -15y=15 Add..................................... -23y = -49 Y = 49/23 Substitution Y = 49/23 solve for x 8x-5(49/23) =5 8x-245/23 =5 x = 5+(245/23) X = (360/23):8 X = 45/23